can you please help me to figure out how to solve this
Caribbean paradise has 35 more suits than pacific paradise. med paradise has 36 fewer than twice the suits of pacific paradise. find the number of suits for each ship total is 519

Found 2 solutions by JulietG, richwmiller:
Answer by JulietG(1812)

C = P+35
M = 2P-36
C + M + P = 519
(P+35) + (2P-36) + P = 519
4P -1 = 519
4P = 520
p = 130

That one is more likely.

Answer by richwmiller(17219)

Perhaps you mean "suites" and not "suits."